OpenStudy (anonymous):

x plus x over 4 equals fifteen, show the steps

OpenStudy (bibby):

\(\huge \frac{x+x}{4}=15\) ? start by simplifying the numerator

OpenStudy (misty1212):

\[x+\frac{x}{4}=15\]??

OpenStudy (bibby):

wait no you probably meant \(\large x+\frac{x}{4}=15\)

OpenStudy (bibby):

get the left side into a common denominator first

OpenStudy (anonymous):

haha yeah sorry. this is for my little brother. I cant figure it out

OpenStudy (bibby):

suprisingly, this is your first math question

OpenStudy (bibby):

anyhow get x into terms of s omething over 4 \(\huge x=\frac{?}{4}\)

OpenStudy (anonymous):

still lost

OpenStudy (camerondoherty):

Hint: Multiply both sides by 4 /.\

OpenStudy (bibby):

you have to get \(x+\frac{x}{4} \) into a single thing of the format \(\large \frac{something}{4}\)

OpenStudy (anonymous):

so x + x/4(4) = 15(4)?

OpenStudy (acxbox22):

no u have to multiply the x by 4 to cuz you are multiplying all of both sides

OpenStudy (camerondoherty):

You multiply both terms on the left side too So \[\Huge x+\frac{ x }{ 4 }=15\] \[\Huge x(4)+\frac{ x }{ 4 }(4)=15(4)\]

OpenStudy (bibby):

you don't have to worry about the right hand side. first step is get the left hand side over a common denominator. so we have \(\huge x+\frac{x}{4}=15\\\huge\frac{4x}{4}+\frac{x}{4}=15\\\huge\frac{4x+x}{4}=15\\\large\frac{5x}{4}=15\)

OpenStudy (anonymous):

We need the final answer to be in the format of, x = 12

OpenStudy (bibby):

well yeah, simplify the last statement

OpenStudy (anonymous):

lol how? i havent done math in years. this is for my little brothers hw

OpenStudy (bibby):

start by multiplying both sides by the denominator, then divide to isolate x further tbh I like cam's method a lot more

OpenStudy (bibby):

\(\large x(4)+\frac{ x }{ 4 }(4)=15(4)\) \(\large 4x+x=60\) \(\large 5x=60\) x=12

OpenStudy (anonymous):

youre my favorite person lol thank you bibby!

OpenStudy (bibby):

hmm wait it might be x=15

OpenStudy (anonymous):

no its 12, we looked it up in the back of the book

OpenStudy (bibby):

damn I fricked up then

OpenStudy (bibby):

wait no I d idn't yay

OpenStudy (bibby):

5x/4=15 5x=60 x=12

OpenStudy (anonymous):

you did it right though lol

OpenStudy (bibby):

I'm so proud of myself